OpenStudy (anonymous):

need help solving this inequality please. x >-7

OpenStudy (anonymous):

...isn't that already solved? You know the values of x. That's about as simple as it gets

OpenStudy (anonymous):

it's a question on an exam prep....

OpenStudy (anonymous):

hm... anything more to it than just "solve..."?

OpenStudy (anonymous):

So what are the values of x that satisfy the relation x > -7

OpenStudy (anonymous):

like choices? Otherwise i think im at a loss

myininaya (myininaya):

(-7,infinity)

OpenStudy (anonymous):

how would I graph it....thats the other part of the question...sorry guys

OpenStudy (anonymous):

\[x \in (-7, \infty)\] Graph the relation? or the solutions?

OpenStudy (anonymous):

on a number line, draw it with an open circle on -7, with an arrow to the right

OpenStudy (anonymous):

Thanks Bwah!

OpenStudy (anonymous):

on a cartesian plane, a vertical line at x=-7, dotted (non inclusive). shade to the right

myininaya (myininaya):

x=-7 is a vertical line so since we have x>-7 and not x>=-7 then you draw a broken vertical line at x=7 and since we have x>-7 shade everything to the right 0f -7

OpenStudy (anonymous):

no prob, glad to help

OpenStudy (anonymous):

how about 3-10x>33

OpenStudy (anonymous):

well, the process is the same as any equation u just have an inequality sign at the end

OpenStudy (anonymous):

Depends on which dimensions he has to include in the graph. In \(R^1\) it's just a number line. In \(R^2\) it is a field comprised of all (x,y) pairs such that \(y\in (-\infty, \infty) \text{ and } x \in (-7, \infty)\) . In \(R^3\) it would be a boxy thing, and so on.

OpenStudy (anonymous):

I think I am lost on the process :)

OpenStudy (anonymous):

How would you solve 3-10x = 33?

OpenStudy (anonymous):

-10x>30 x<-3 (if i recall correctly, u switch the sign if u divide or multiply by a negative number)

OpenStudy (anonymous):

subtract 3 from both sides?

OpenStudy (anonymous):

Indeed. You can do the same thing to inequalities.

OpenStudy (anonymous):

Then what would you do?

OpenStudy (anonymous):

divide by 10......I think I am seeing the light :)

OpenStudy (anonymous):

Yep. The only rule is that if you divide or multiply by a negative you have to change the direction of the inequality. \(-x < 5 \implies x > 5\)

OpenStudy (anonymous):

Ack

OpenStudy (anonymous):

\[-x < 5 \implies x > -5\] rather

OpenStudy (anonymous):

Thanks so very much!

myininaya (myininaya):

I don't know if you have ever though of like if we have 5>2 then and we multiply both sides by negative one or divide by negative one then it becomes -5<-2 since -2 is bigger than -5

myininaya (myininaya):

any any negative number